OpenStudy (anonymous):

To find the height of a tall tree, a surveyor moves 140 feet away from the base of the tree and then, with a transit 4 feet tall, measures the angle of elevation to the top of the tree to be 53°. What is the height of the tree? Round your answer to the nearest foot. A. 190 ft B. 186 ft C. 105 ft D. 109 ft

OpenStudy (timaashorty):

\[ 4 + 140 \times \tan(53) = 186.7862\]

OpenStudy (anonymous):

tan 53 = (h-4)/ 140 Solve for h